If a local pizza parlor has three different crusts and 15 different toppings to choose from, in how many different ways could you order a four-toping pizza? (One crust, four topping).
Answer #1

number of ways=N(choose one crust from 3 and choose 4 topping from 15)

=(3C1)*(15C4) =3*1365=4095